# Machine Learning and First-Principles Predictions of Materials with Low Lattice Thermal Conductivity

### v1

@@ I Introduction
The TE performance can be quantified by the figure of merit value $ZT=\frac{S^{2}\sigma}{\kappa T}$, where $S$ is the Seebeck coefficient (the induced voltage in response to a temperature gradient), $\sigma$ is the electrical conductivity, $\kappa$ is the thermal conductivity, and $T$ is the temperature. One approach to enhance $ZT$ is by increasing the power factor ($S^{2}\sigma$), via band engineering of carrier concentration and mobility, among other factors [9, 10, 11, 12, 13, 14, 15]. Another approach is to find materials with low thermal conductivity [16, 17, 18].
@@ II.2 First-Principles Calculation
Figure 1: (a) Histogram for the 1900 training samples of lattice thermal conductivity ($\kappa_{L}$) selected from the TE Design Lab [47]. The distribution spans nearly five orders of magnitude and is plotted on a logarithmic scale. (b) False-color intensity plot showing the frequency of each element in the 1900 training dataset. Elements not present in the list are shown in gray. The figure was created using the open-source software Periodic Trend Plotter [64].
@@ III RESULTS AND DISCUSSION
Figure 1(a) shows the distribution of $\kappa_{L}$ for the 1900 compounds in our training dataset from the TE Design Lab [47]. Since the range of the distribution spans nearly 5 orders of magnitude, it is plotted on a logarithmic scale. Eventually, ML models are trained to predict $\log(\kappa_{L})$. For the accuracy and generalizability of our ML models, we ensure that our dataset is diverse in chemical composition (from unary to quinary compounds) and crystal structure (with 140 different space groups). In particular, among the 1900 samples, 7 are unary, 418 are binary, 1143 are ternary, 328 are quaternary, and 4 are quinary. These compounds contain 61 different atomic elements; the frequency each element appears in the compound list is represented by the false-color intensity plot in Figure 1(b). Gray color means that the element is not present.
@@
We next apply the ML models to predict materials with low $\kappa_{L}$. Recently, Zintl-phase compounds have attracted significant attention due to their strong anharmonic properties, which could lead to low lattice thermal conductivities [70, 71, 72, 73, 74, 75, 76, 77]. The Zintl phase refers to compounds formed by alkali metals (group I) or alkaline-earth metals (group II) combined with $p$-block metals or metalloids (from groups III-VI). Other recent studies have also shown that Zintl-phase compounds can achieve ultra-low $\kappa_{L}$ by introducing a heavy element, cadmium (Cd) [78, 20]. For these reasons, we focus on applying our ML models to Cd-based Zintl-phase materials. Specifically, we consider A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) with orthorhombic symmetry and space group Ama2 (No. 40) [79]; Figure 3(a) shows the corresponding crystal structure for K2CdPb. As seen in Table I, the $\kappa_{L}$ values predicted by our ML models for the nine compositions A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) range from $0.69$ to $0.95$ W/mK, indicating that all these compounds are potential low-$\kappa_{L}$ materials.
Figure 3: (a) Primitive-cell crystal structure of K2CdPb, with orthorhombic symmetry and space group Ama2 (No. 40). (b) Lattice thermal conductivity $\kappa_{L}$ of K2CdPb as a function of the neighbor interaction cutoff distance in the temperature range of $300-1200$ K. (c) $\kappa_{L}$ of K2CdPb as a function of temperature for different cutoff distances. (d) $\kappa_{L}$ computed with a cutoff distance of 4 Å for nine different Cd-compounds.
To validate the ML predictions, we further perform first-principles calculations to directly compute $\kappa_{L}$ and other thermoelectric properties for the nine compounds under study. We first focus on K2CdPb [79], whose primitive cell structure is shown in Figure 3(a). To ensure accurate calculation of $\kappa_{L}$, we conduct a convergence test for the neighbor interaction cutoff distance. Figures 3(b) and 3(c) show the convergence tests for K2CdPb as functions of temperature and cutoff distance. Notably, the $\kappa_{L}$ computed with a cutoff of 4 Å is very close to the result obtained with a 5 Å cutoff. Therefore, for both accuracy and efficiency considerations, we have adopted a cutoff distance of 4 Å for the other compounds. Figure 3(d) shows the computed $\kappa_{L}$ as a function of temperature for the nine compounds A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge). Near room temperature, all compounds exhibit a $\kappa_{L}$ below 1.0 W/mK, in very good agreement with our ML predictions. As the temperature increases, $\kappa_{L}$ further decreases as more phonons are excited and cause additional phonon scattering, leading to a reduction in $\kappa_{L}$. These results reveal that the nine compounds under study are all low-$\kappa_{L}$ materials for potential thermal management and insulation applications.
@@
Before discussing other thermoelectric properties, we address the small discrepancies between the ML and DFT results in Table I. First, we note that the ML models were trained to predict $\log(\kappa_{L})$ rather than $\kappa_{L}$ itself. Therefore, a small error in the logarithmic value can be amplified in the actual value. Second, the ML prediction in Table I is generally slightly larger than the DFT calculation. One reason for this discrepancy is likely due to the training data distribution. Specifically, while we are interested in discovering materials with low thermal conductivity (i.e., $\log(\kappa_{L})\leq 0$), most of the training data in Figure 1(a) exhibit $\log(\kappa_{L})\geq 0$. Therefore, one could potentially enhance the ML prediction accuracy by training a weighted ML model. Indeed, when we train additional ML models by weighting the training samples with $\log(\kappa_{L})\leq 0$ by a factor of 5 to 10, the predicted $\kappa_{L}$ values become smaller and align more closely with the DFT results, as seen in Table I. Notably, the performance of the ML model with a weight factor of 10 is not better than that with a weight factor of 5. Therefore, the sample weight factor has an optimal range and cannot be increased indefinitely. Finally, we note that some ML predictions from the weighted ML models remain larger than the DFT results (e.g., for the sodium compounds in Table I). This discrepancy may be attributed to the fact that tree models only interpolate so cannot predict values beyond the range of the training dataset. The apparent, albeit small, differences between the ML and DFT predictions are likely associated with the above factors.
@@
Finally, since low-$\kappa$ materials can be good candidates for thermoelectrics, we also compute their figure of merit, $ZT=\frac{S^{2}\sigma}{\kappa T}$, where the thermal conductivity $\kappa=\kappa_{e}+\kappa_{L}$ includes both electronic and lattice contributions. Figures 5(a)-(c) show the $ZT$ values respectively for K2CdPb, K2CdSn, and K2CdGe, as functions of carrier concentration $n$ (in log scale) and temperature $T$. In all three compounds, the $ZT$ values can exceed 1.0. In particular, K2CdPb exhibits a peak $ZT$ value of $\sim 1.4$ at $n=2\times 10^{20}$ cm-3 and $T=400$ K, making it a promising low-temperature thermoelectric material. In contrast, K2CdSn and K2CdGe show peak $ZT$ values of $\sim 1.1$ near $n=9\times 10^{20}$ cm-3 and $T=900$ K, and they are more suitable for thermoelectric applications at higher temperatures. For the other compounds based on sodium and lithium listed in Table I, the $ZT$ values are less than 1.0, making them unsuitable for practical thermoelectric applications. However, they could still be potential candidates for thermal insulation materials.

### v2

@@ I Introduction
The TE performance can be quantified by the figure of merit value $ZT=S^{2}\sigma T/\kappa$, where $S$ is the Seebeck coefficient (the induced voltage in response to a temperature gradient), $\sigma$ is the electrical conductivity, $T$ is the temperature, and $\kappa$ is the thermal conductivity. One approach to enhance $ZT$ is by increasing the power factor ($S^{2}\sigma$), via band engineering of carrier concentration and mobility, among other factors [9, 10, 11, 12, 13, 14, 15]. Another approach is to find materials with low thermal conductivity [16, 17, 18].
@@ II.1 Machine Learning Simulation
We note that several features, such as average atomic mass and volume (which is related to atomic radius), are relevant parameters for estimating $\kappa_{L}$ in known empirical formulas [50, 51]. There, $\kappa_{L}$ is also expected to be proportional to the mean sound velocity $v_{m}$ (or the Debye temperature $\Theta_{D}$) cubed. It was shown that ML models can accurately predict $v_{m}$ and $\Theta_{D}$ using features simply derived from chemical compositions and crystal symmetry [52]. Therefore, it is anticipated that ML models trained here with the 61 features under study can perform well in predicting $\kappa_{L}$ [49].
@@ III Results and Discussion
Figure 1: (a) Histogram for the 1900 training samples of lattice thermal conductivity ($\kappa_{L}$) selected from the TE Design Lab [47]. The distribution spans nearly five orders of magnitude and is plotted on a base-10 logarithmic scale. (b) False-color intensity plot showing the frequency of each element in the 1900 training dataset. Elements not present in the list are shown in gray. The figure was created using the open-source software Periodic Trend Plotter [67].
Figure 1(a) shows the distribution of $\kappa_{L}$ for the 1900 compounds in our training dataset from the TE Design Lab [47]. Since the range of the distribution spans nearly 5 orders of magnitude, it is plotted on a base-10 logarithmic scale. Eventually, ML models are trained to predict $\log(\kappa_{L})$. For the accuracy and generalizability of our ML models, we ensure that our dataset is diverse in chemical composition (from unary to quinary compounds) and crystal structure (with 140 different space groups). In particular, among the 1900 samples, 7 are unary, 418 are binary, 1143 are ternary, 328 are quaternary, and 4 are quinary. These compounds contain 61 different atomic elements; the frequency each element appears in the compound list is represented by the false-color intensity plot in Figure 1(b). Gray color means that the element is not present.
@@
We next apply the ML models to predict materials with low $\kappa_{L}$. Recently, Zintl-phase compounds have attracted significant attention due to their strong anharmonic properties, which could lead to low lattice thermal conductivities [71, 72, 73, 74, 75, 76, 77, 78]. The Zintl phase refers to compounds formed by alkali metals (group I) or alkaline-earth metals (group II) combined with $p$-block metals or metalloids (from groups III-VI). Other recent studies have also shown that Zintl-phase compounds can achieve ultra-low $\kappa_{L}$ by introducing a heavy element, cadmium (Cd) [79, 20]. For these reasons, we focus on applying our ML models to Cd-based Zintl-phase materials. Specifically, we consider A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) with orthorhombic symmetry and space group Ama2 (No. 40) [80]; Figure 3(a) shows the corresponding crystal structure for K2CdPb. As seen in Table 1, the $\kappa_{L}$ values predicted by our ML models for the nine compositions A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge) range from $0.69$ to $0.95$ W/mK, indicating that all these compounds are potential low-$\kappa_{L}$ materials.
Figure 3: (a) Primitive-cell crystal structure of K2CdPb, with orthorhombic symmetry and space group Ama2 (No. 40). (a) Lattice thermal conductivity $\kappa_{L}$ of K2CdPb as a function of the neighbor interaction cutoff distance in the temperature range of $300-1200$ K. (a) $\kappa_{L}$ of K2CdPb as a function of temperature for different cutoff distances. (d) $\kappa_{L}$ computed with a cutoff distance of 4 Å for nine different Cd-compounds.
@@
To validate the ML predictions, we further perform first-principles calculations to directly compute $\kappa_{L}$ and other thermoelectric properties for the nine compounds under study. We first focus on K2CdPb [80], whose primitive cell structure is shown in Figure 3(a). To ensure accurate calculation of $\kappa_{L}$, we conduct a convergence test for the neighbor interaction cutoff distance. Figures 3(b) and 3(c) show the convergence tests for K2CdPb as functions of temperature and cutoff distance. Notably, the $\kappa_{L}$ computed with a cutoff of 4 Å is very close to the result obtained with a 5 Å cutoff. Therefore, for both accuracy and efficiency considerations, we have adopted a cutoff distance of 4 Å for the other compounds. Figure 3(d) shows the computed $\kappa_{L}$ as a function of temperature for the nine compounds A2CdX (A = Li, Na, and K; X = Pb, Sn, and Ge). Near room temperature, all compounds exhibit a $\kappa_{L}$ below 1.0 W/mK, in very good agreement with our ML predictions. As the temperature increases, $\kappa_{L}$ further decreases as more phonons are excited and cause additional phonon scattering, leading to a reduction in $\kappa_{L}$. These results reveal that the nine compounds under study are all low-$\kappa_{L}$ materials for potential thermal management and insulation applications. We further note that our reported $\kappa_{L}\sim 0.3$ W/mK for K2CdPb and K2CdSn is consistent with previous DFT studies [80]. Meanwhile, we also find that K2CdGe exhibits a comparable theoretical $\kappa_{L}\sim 0.3$ W/mK. This result is not surprising, given the chemical similarity of the Pb, Sn, and Ge elements. Likewise, for the Li- and Na-based compounds explored here, while our DFT calculations may underestimate their $\kappa_{L}$ values, they are also anticipated to be low lattice thermal conductivity materials, based on their chemical similarity to the K-based compounds. In fact, depending on carrier concentration and the underlying temperature, materials with ultra-low $\kappa_{L}\sim 0.1$ W/mK or lower have been reported in the literature [81, 82, 83, 84]. It would be an important future task to verify our predictions both theoretically (e.g., with different DFT functionals and supercell sizes) and experimentally, through potential synthesis and characterization of the proposed materials.
Before discussing other thermoelectric properties, we address the small discrepancies between the ML and DFT results in Table 1. First, we note that the ML models were trained to predict $\log(\kappa_{L})$ rather than $\kappa_{L}$ itself. Therefore, a small error in the logarithmic value can be amplified in the actual value. Second, the ML prediction in Table I is generally slightly larger than the DFT calculation. One reason for this discrepancy is likely due to the training data distribution. Specifically, while we are interested in discovering materials with low thermal conductivity (i.e., $\log(\kappa_{L})\leq 0$), most of the training data in Figure 1(a) exhibit $\log(\kappa_{L})\geq 0$. Therefore, one could potentially enhance the ML prediction accuracy by training a weighted ML model. Indeed, when we train additional ML models by weighting the training samples with $\log(\kappa_{L})\leq 0$ by a factor of 5 to 10, the predicted $\kappa_{L}$ values become smaller and align more closely with the DFT results, as seen in Table I. Notably, the performance of the ML model with a weight factor of 10 is not better than that with a weight factor of 5. Therefore, the sample weight factor has an optimal range and cannot be increased indefinitely. Finally, we note that some ML predictions from the weighted ML models remain larger than the DFT results (e.g., for the sodium compounds in Table I). This discrepancy may be attributed to the fact that tree models only interpolate so cannot predict values beyond the range of the training dataset. The apparent, albeit small, differences between the ML and DFT predictions are likely associated with the above factors.
@@
Finally, since low-$\kappa$ materials can be good candidates for thermoelectrics, we also compute their figure of merit, $ZT=S^{2}\sigma T/\kappa$, where the thermal conductivity $\kappa=\kappa_{e}+\kappa_{L}$ includes both electronic and lattice contributions. Figures 5(a)-(c) show the $ZT$ values respectively for K2CdPb, K2CdSn, and K2CdGe, as functions of carrier concentration $n$ (in log scale) and temperature $T$. In all three compounds, the $ZT$ values can exceed 1.0. As an example to estimate the $ZT$ value, for K2CdPb at $T=400$ K and $n=2\times 10^{20}$ cm-3, the relevant parameters from our calculations are $S\sim 1.8\times 10^{-4}$ V/K, $\sigma\sim 0.1\times 10^{5}$ 1/$\Omega$m, and $\kappa=\kappa_{e}+\kappa_{L}\sim 1.0$ W/mK. Together, these values lead to a figure of merit $ZT=S^{2}\sigma T/\kappa(\sim 3.24\times 10^{-8}\times 0.1\times 10^{5}\times 400/1.0)\sim 1.3-1.4$ for K2CdPb, making it a promising low-temperature thermoelectric material. In contrast, K2CdSn and K2CdGe show peak $ZT$ values of $\sim 1.1$ near $n=9\times 10^{20}$ cm-3 and $T=900$ K, and they are more suitable for thermoelectric applications at higher temperatures. For the other compounds based on sodium and lithium listed in Table I, the $ZT$ values are less than 1.0, making them unsuitable for practical thermoelectric applications. However, they could still be potential candidates for thermal insulation materials.